OpenStudy (anonymous):

Prove that :If a,b belongs to Z then (a^2 + b^2)/(1 + ab) is a perfect square.

OpenStudy (anonymous):

I know the answer......I just want to confirm it.

OpenStudy (anonymous):

a = 2, b =1 \[\frac{4 + 1}{1 +2} = \frac{5}{3}\]

OpenStudy (anonymous):

Or, I didn't understand your question? I don't think you will ever get a perfect square through that equation.

OpenStudy (anonymous):

Try b = 0 you get one . Now how do I prove it?

OpenStudy (anonymous):

This is an IMO Problem.

OpenStudy (anonymous):

But I think that's it, it's only valid when b=0. Do you have any other example?

OpenStudy (anonymous):

\[\frac{a^2+b^2}{1+ab} = q^2\] \[\sqrt{\frac{a^2+b^2}{1+ab}} = q\]Now we need some manipulation here.